the ratio of length and breadth of rectangle is 5:3.Find l and b if perimeter is 256
Answer by TimothyLamb(4379)

---
x = length
y = breadth
---
x/y = 5/3
3x = 5y
---
2x + 2y = 256
---
linear system:
3x = 5y
2x + 2y = 256
---
put the system of linear equations into standard form:
---
3x - 5y = 0
2x + 2y = 256
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = length = 80
y = breadth = 48
---
system matrix
3 -5
2 2
---
inverse of system matrix
0.125 0.3125
-0.125 0.1875
---
determinant of system matrix = 16
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php